How much solute is required to make 250mL of a 1 molar solution of KCl? Given K= 39.1g Cl=35.45g
Answer #1

Molarity = (W /MW) (1000 / V in mL )

Molarity = 1.0 M

W = ?

MW = 74.55 g /mol

V = 250 mL

1.0 = (W / 74.55) (1000/250)

W / 74.55 = 0.25

W = 18.64 g

mass of solute required = 18.64 g